OpenStudy (love_to_love_you):

PLEASE HELP WILL FAN AND MEDAL Write a polynomial function with the given zeros. Show all work. a. x= -1, 0, 2 b. x= -2, 1, 4

OpenStudy (love_to_love_you):

@campbell_st

OpenStudy (campbell_st):

ok... so if you know a zero if x = -1 is a zero or root, then (x + 1) is a factor x = 2 is zero or root then (x -2) is a factor x = 0 is a zero then x is a factor so the polynomial is P(x) =x(x-2)(x+1) distribute for the standard form (b) use the same reasoning to identify the factors. hope it helps

OpenStudy (love_to_love_you):

if x = -2 is a zero or root, then (x + 2) is a factor x = 1 is zero or root then (x - 1) is a factor x = 4 ? what do i do here?

OpenStudy (love_to_love_you):

@campbell_st

OpenStudy (love_to_love_you):

@ganeshie8

OpenStudy (love_to_love_you):

@Loser66

OpenStudy (love_to_love_you):

@AlexandervonHumboldt2

OpenStudy (campbell_st):

if x = 4 is a zero then (x - 4) is a factor so P(x) = (x +2)(x-1)(x-4) just distribute

OpenStudy (love_to_love_you):

Thank you!